Circular No. 9023

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2008 S 1 C. C. Porco, on behalf of the Cassini Imaging Science Team, reports

the discovery of S/2008 S 1, an object orbiting within Saturn’s G ring in the region between Saturn X/XI (Janus/Epimetheus) and Saturn I (Mimas), in two images of the G ring obtained with Cassini’s narrow-angle camera taken on 2008 Aug. 15 and separated by ≈ 30 min. Subsequent observations, and a search of earlier images based on a preliminary orbit fit, have provided additional detections of this object, yielding 21 total detections during 2007 June 15–2009 Feb. 20. A high-precision orbit fit based on a numerical integration of the equations of motion shows that the object is being influenced by a 7:6 mean-motion resonance with Mimas; averaged orbital elements for S/2008 S 1: a = 167500 km; e = 0.0002, i = 0^{o}.001, revolution period = 0.80812 day. The integrated orbit shows that the resonance with Mimas causes a long-period (∼ 4 yr) oscillation of ∼ 4 km in the semi-major axis of S/2008 S 1, with a corresponding oscillation of a few degrees in its mean longitude. S/2008 S 1 is located within a bright arc in the G ring, so this body likely represents a significant source of the small particles found in this ring. If S/2008 S 1 has the same disk-integrated reflectivity as Saturn XXXIII (Pallene), preliminary estimates of the object’s total brightness suggest a physical radius for S/2008 S 1 of ∼ 250 m.

COMET C/2007 N3 (LULIN) D. E. Harker, University of California, San Diego; and C. E. Wood-

ward, University of Minnesota, report on spectrophotometric observations of comet C/2007 N3 using the NASA Infrared Telescope Facilty (IRTF) telescope (+ MIRSI) on Feb. 22.6 UT. The coma was spatially extended at mid-infrared wavelengths, with a circularly symmetric extent of ≈ 6′′ in diameter (3σ surface brightness). Photometry in a 3′′.18-diameter circular aperture, centered on the peak in the coma surface brightness coincident with the nucleus position, yields fluxes of within the narrow-band IRTF comet-silicate filters of [8.7 μm] = 3.58 ± 0.06 Jy, [9.8 μm] = 4.98 ± 0.09 Jy, [10.6 μm] = 5.99 ± 0.03 Jy, [11.6 μm] = 6.82 ± 0.04 Jy, and [12.3 μm] = 7.54 ± 0.06 Jy. A Planck-function fit to the photometry yields a color temperature of 267 ± 10 K, which is 14 percent higher than that of a blackbody at an equivalent heliocentric distance (1.38 AU). Both the photometric spectral energy distribution and a low-resolution 10-μm spectrum indicate that C/2007 N3 has no silicate emission feature, implying a deficiency of small, submicron-sized silicate grains in the coma.

2009 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green